OpenStudy (anonymous):

ABCDE and FGHJE are similar pentagons. If the perimeter of FGHJE is 7.3, what is the perimeter of ABCDE? 14 14.6 15 14.5

OpenStudy (phi):

similar pentagons means that one pentagon is a "multiple" of the other... in other words, you can multiply the perimeter of one by some number, and get the perimeter of the other

OpenStudy (anonymous):

so is B or D

OpenStudy (phi):

You need more info to answer the question. For example, do they show the length of a side ?

OpenStudy (phi):

I assume ABCDE is bigger than the other one, but how much bigger ?

OpenStudy (anonymous):

A to be = 3.5 and F to G= 1.75

OpenStudy (phi):

to do this problem, you need to find the number that when you multiply it times 1.75 you get 3.5 any idea how to find that number ?

OpenStudy (anonymous):

no

OpenStudy (phi):

think like this some number (call it "k") times 1.75 is 3.5 we write 1.75 k = 3.5 this is the same kind of problem as 2*k = 4 what is k? we find it by dividing 4 by 2 k = 4/2 for your problem, do the same thing k = 3.5/1.75

OpenStudy (phi):

can you find the number ?

OpenStudy (phi):

you could use google, type 3.5/1.75= into google

OpenStudy (anonymous):

is 2

OpenStudy (phi):

that means the bigger pentagon is 2 times bigger than the smaller one. Specifically, it means its perimeter is 2 times bigger than the perimeter of the smaller pentagon

OpenStudy (anonymous):

oh okay. I guess i try to see and THANK YOU SO MUCH.

OpenStudy (phi):

to get the answer you do 2*7.3 you can use google (but you should be able to do that multiply in your head)

OpenStudy (anonymous):

IS 14.6 IS B I ALREADY FOUND IT THANK U SO MUCH.....